### v1

# High-dimensional Autoregressive Modeling for Time Series with Hierarchical Structures

Abstract
High-dimensional time series often exhibit hierarchical structures represented by tensors, while statistical methodologies that can effectively exploit the structural information remain limited. We propose a supervised factor modeling framework that accommodates general hierarchical structures by extracting low-dimensional features sequentially in the mode orders that respect the hierarchical structure. Our method can select a small collection of such orders to allow for impurities in the hierarchical structures, yielding interpretable loading matrices that preserve the hierarchical relationships. A practical estimation procedure is proposed, with a hyperparameter selection scheme that identifies a parsimonious set of action orders and interim ranks, thereby revealing the possibly latent hierarchical structures. Theoretically, non‑asymptotic error bounds are derived for the proposed estimators in both regression and autoregressive settings. An application to the IPIP‑NEO‑120 personality panel illustrates superior forecasting performance and clearer structural interpretation compared with existing methods based on tensor decompositions and hierarchical factor analysis.
Keywords: Autoregressive model, Hierarchical tensor data, High-dimensional time series, Non-asymptotic properties, Supervised factor modeling
1 Introduction
High-dimensional time series abound in modern scientific and econometric studies (wang2019factor; chang2024autocovariance; chen2025estimating), which often exhibit multiway (chen2022analysis; chen2022factor; wang2024high) and partially hierarchical organizations (johnson2014measuring; gao2023divide). The data are hierarchical when they contain multiple modes representing the layers of a nested structure obtained from several groups, subgroups, and items. In many applications, however, the hierarchical structure is only partial as some modes lie outside this hierarchy or even form another hierarchy. For balanced data which admits a tensor representation, methods that can adapt to such partially hierarchical tensor time series remain limited, as existing tensor-based methods or high-dimensional dynamic factor models typically assume a fully juxtaposing or hierarchical organization (montgomery2006design; stock2011dynamic; brunner2012tutorial; yuan2023two). In practice, real datasets frequently combine both: hierarchical components accompanied by cross-cutting covariates. Developing a framework that can efficiently summarize, predict, and statistically interpret such partially hierarchical structures with complex inter-mode dependencies is therefore of both theoretical and practical importance.
A representative example is the International Personality Item Pool (IPIP) NEO-120 personality inventory (johnson2014measuring; qiao2025exploratory), which contains 120 questions assessing five broad personality traits, each with 24 questions. The data are collected over time from participants of different ages and genders, resulting in a three-way tensor time series with dimensions $12\penalty 10000\ (gender\penalty 10000\ and\penalty 10000\ age\penalty 10000\ groups)\times 5\penalty 10000\ (personalities)\times 24\penalty 10000\ (questions)$. The last two dimensions reflect a hierarchical organization, while the first does not belong to this hierarchy. Given the dataset’s scale ($12\times 5\times 24=1440$ dimensions) and the limited number of time points (approximately 100 monthly observations), there is a need for an efficient and statistically interpretable model that can extract latent features from the data for prediction of future observations. The extraction process which reduces the dimensionality in a supervised manner should respect the data’s partially hierarchical structure, when both latent feature detection and trend analysis are of sociological interest.
Natural strategies to handle data of such high dimensionality include imposing sparsity or low-rankness on the model, so that a more tractable low-dimensional structure emerges. Sparsity-inducing methods (basu2015regularized; raskutti2019convex; bertsimas2020sparse) have been widely used in high-dimensional time series analysis with its well-known prediction ability (wilms2023sparse; zheng2025interpretable). However, these methods may lack structural interpretability unless the underlying sparsity pattern aligns with a meaningful structure, as is common in genetics and neuroscience (zhou2013tensor; bi2022modeling; wu2023sparse). In IPIP-NEO-120, for instance, the responses within each personality trait are highly correlated due to the experimental design, and a sparse model may fail to capture this inherent structure.
In contrast, low-rank modeling compresses the data to capture latent structure via orthogonal loading matrices, leading to a more interpretable framework of factor modeling and analysis. Supervised Factor Modeling (SFM) (huang2025supervised) is a prominent low-rank approach that uses loading matrices to extract features (or factors) from both predictors and responses, followed by a multivariate regression on these features. Specifically for hierarchical data, literature is copious with methods that extract features tailored to the hierarchical construct, such as higher-order factor models (yung1999relationship), group factor models (klami2014group), hierarchical factor models (schmid1957development; qiao2025exact). These methods either extracts factors recursively along the hierarchy, or directly from each of its levels. However, they all assign different loading matrices to different lower-level modes under the same higher-level mode, so the number of parameters exceed the dimensions of the data. These approaches are thus only feasible in the unsupervised context, where the proportion of explained variance, rather than prediction performance, is the primary concern. When otherwise applied in the supervised context, only low dimensional settings are allowed.
As traditional low-rank methods for unsupervised factor analysis of hierarchical constructs are unsuitable for high-dimensional estimation problem, SFM methods based on tensor decomposition (huang2025supervised) are preferred given their parsimony. Among these, Tucker and canonical polyadic (CP) decompositions are widely used (kolda2009tensor). Tucker decomposition (tucker1966some) is notable for its numerical stability and interpretable orthogonal loadings, which align well with supervised factor modeling (wang2024high). However, its model complexity grows exponentially with tensor order, limiting its efficiency in high-order settings (oseledets2011tensor). This issue is exacerbated in tensor autoregressive models, where coefficient tensors often have high orders. In the IPIP-NEO-120 example, a lag-one tensor autoregressive model would involve a sixth-order coefficient tensor, and estimation based on Tucker decomposition would be computationally less efficient. Additionally, Tucker loadings are confined on Kronecker product spaces, restricting its applicability to data with complex or hierarchical structures, as discussed in Sections 2.1 and 2.2. CP decomposition (harshman1970foundations) offers a more efficient and flexible alternative (lock2018tensor), with model complexity scaling linearly with tensor order and dimensions. However, CP loadings are complicated, constraining its interpretation in unsupervised settings as a PCA surrogate (han2024cp; han2022tensor), and CP-based SFM (cai2025efficient) lacks clear interpretation for structured data.
Table 1: Comparison of different methods for supervised factor modeling. For simplicity, we assume equal dimension $p$ for each mode of the data, and equal number of features, $r$, extracted from each mode, and denote $d$ as the order of the coefficient tensor.
Models SFM based on Tensor Decomp. Factor Models for SFM
Ours CP Tucker $T$ GFM HFM HOFM
Number of parameters: $O(\cdot)$ $pr^{2}d$ prdprd $prd+r^{d}$ $pr^{2}d$ $rp^{d}$ $rp^{d}$ $rp^{d}$
Interpretation: Hierarchical $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Interpretation: Crossed $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
Meanwhile, tensor regression based on tensor train (TT) decomposition (oseledets2011tensor) of the coefficient tensor has been proposed with its efficiency highlighted for tensors with high orders (liu2020low). Under the SFM structure, qin2025computational and si2024efficient adopt a pair of TT-styled loadings for the predictor and repsonse, and their sequential structure can capture hierarchical relationships among modes with reduced number of parameters. Specifically, each loading matrix is a product of several component matrices, with each component extracting features from one mode to which the previous mode imparts extracted features. The TT-based model is hence interpretable in the regression tasks where both the predictor and response exhibit a fully hierarchical structure (zhou2022optimal; si2024efficient). However, its strict sequential structure inherited from TT decomposition limits its applicability to data with partially hierarchical structures, as in the IPIP-NEO-120 example. Another major limitation is that the order of the component matrices in TT loadings is fixed to be forward for the predictor and backward for the response. This fixed construction hinders its application to autoregressive time series models, where the predictor and response share basically the same structure. Permuting the order of the predictor’s modes for a forced structural alignment in turn deteriorates the autoregressive model into a general stochastic regression model, complicating the temporal dependence analysis. Table 1 provides a summary of the reviewed methods under the SFM framework, comparing their number of parameters and interpretability for hierarchical and crossed (non-hierarchical) structures.
In view of the aforementioned challenges, we propose a novel tensor-based supervised factor modeling framework that effectively extracts features from partially hierarchical tensor time series, while preserving the hierarchical structure information. Our model combines admissible mode orders of the partially hierarchical structure instrinc to the data, analogous to a mixture model. Then, it extracts features sequentially, thereby capturing complex inter-mode dependencies with cross-cutting covariates obscuring any exact hierarchical structure. The proposed method is computationally efficient, as information summarization and its upward impartation to the higher-level layer bear a shared low-rank structure for the lower-level layer. This parsimony enables both accurate estimation with sample size and statistical interpretability inherited from SFM framework. Specifically, our contributions are summarized as follows:
  • We are, to the best of our knowledge, the first to bridge the methodological gap between fully hierarchical structures and non-hierarchical (juxtaposing) structures. Our proposed novel tensor-based SFM method accommodates to partially hierarchical structures in both regression and autoregression tasks, with theoretical guarantees evidenced by both simulation and empirical examples.
  • Our proposed method offers a flexible and unified way to detect from data principal and lesser hierarchical structures forming a mixture, enabled by our hyperparameter selection algorithm, when the true structure is obscured.
  • The construction in our proposed method can reveal an algebraic-graphical duality of the hierarchical structures, enshrined in an algebraic proposition, which may be of independent interest in the study of tensor decomposition and hierarchical models.
The rest of the paper is organized as follows. In Section 2, we introduce the general supervised factor modeling framework, and then present our proposed method for partially hierarchical data, followed by an interpretable tensor autoregressive model. Section 3 provides the theoretical analysis of our method and Section 4 describes its algorithms along with hyperparameter selection. In Section 5, we conduct extensive simulation studies to evaluate the performance of our method against existing alternatives. In Section 6, we apply our method to the IPIP-NEO-120 dataset for analysis. Finally, we conclude with a discussion in Section 7. All technical proofs and additional numerical results are relegated to the Supplementary Material.
Notations and preliminaries.
Throughout the paper, we denote vectors by boldface small letters, e.g., $\boldsymbol{x}$; matrices by boldface capital letters, e.g., $\boldsymbol{X}$; tensors by boldface Euler capital letters, e.g., $\mathscr{X}$. For a $d$-th order tensor $\mbox{$\mathscr{X}$}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{d}}$, denote by $\operatorname*{vec}(\mbox{$\mathscr{X}$})$, $\mbox{$\mathscr{X}$}_{(s)}$, and $\mbox{$\mathscr{X}$}]_{s}$ its vectorization, mode matricization, and sequential matricization at mode $s$ with $1\leq s\leq d$, respectively. The Frobenius norm of $\mathscr{X}$ is defined as $\left\lVert\mbox{$\mathscr{X}$}\right\rVert_{\mathrm{F}}=\sqrt{\langle\mbox{$\mathscr{X}$},\mbox{$\mathscr{X}$}\rangle}$, where $\langle\cdot,\cdot\rangle$ is the inner product. Moreover, the $(i_{1}i_{2}\ldots i_{d})$-th element of $\mathscr{X}$ is denoted by $\mbox{$\mathscr{X}$}_{i_{1}i_{2}\ldots i_{d}}$, for $i_{k}=1,2,\ldots,p_{k}$, $k=1,2,\ldots,d$. We denote $\times_{s}$ the $s$-th mode matrix product, and $\otimes$ denotes the Kronecker product. For a generic matrix $\boldsymbol{X}$, we denote by $\boldsymbol{X}^{\top}$, $\left\lVert\boldsymbol{X}\right\rVert_{\mathrm{F}}$, $\left\lVert\boldsymbol{X}\right\rVert_{2}$, and $\operatorname*{vec}(\boldsymbol{X})$ its transpose, Frobenius norm, spectral norm, and a long vector obtained by stacking all its columns, respectively. If $\boldsymbol{X}$ is further a square matrix, then denote its minimum and maximum eigenvalue by $\lambda_{\min}(\boldsymbol{X})$ and $\lambda_{\max}(\boldsymbol{X})$, respectively. Denote the orthogonal set as $\mathbb{O}^{p\times r}=\{\boldsymbol{X}\in\mathbb{R}^{p\times r}:\boldsymbol{X}^{\top}\boldsymbol{X}=\boldsymbol{I}_{r}\}$ for $p\geq r$, where $\boldsymbol{I}_{r}\in\mathbb{R}^{r\times r}$ is the identity matrix. With a little abuse of terminology, a rectangular matrix is said to be orthogonal if it belongs to $\mathbb{O}^{p\times r}$. For two real-valued sequences $x_{k}$ and $y_{k}$, $x_{k}\gtrsim y_{k}$ if there exists a $C>0$ such that $x_{k}\geq Cy_{k}$ for all $k$. In addition, we write $x_{k}\asymp y_{k}$ if $x_{k}\gtrsim y_{k}$ and $y_{k}\gtrsim x_{k}$.
2 Model settings
2.1 Tucker tensor regression and supervised factor modeling
We begin with a general regression model involving tensor-valued responses and predictors:
[…]
$\mbox{$\mathscr{A}$}=\mbox{$\mathscr{G}$}\times_{1}\boldsymbol{U}_{1}\times_{2}\cdots\times_{N}\boldsymbol{U}_{N}\times_{N+1}\boldsymbol{V}_{1}\times_{N+2}\cdots\times_{N+M}\boldsymbol{V}_{M},$ (2.2)
where $\mbox{$\mathscr{G}$}\in\mathbb{R}^{s_{1}\times\cdots\times s_{N}\times r_{1}\times\cdots\times r_{M}}$ is the core tensor, and $\boldsymbol{U}_{n}\in\mathbb{R}^{q_{n}\times s_{n}}$ with $1\leq n\leq N$ and $\boldsymbol{V}_{m}\in\mathbb{R}^{p_{m}\times r_{m}}$ with $1\leq m\leq M$ are factor matrices. Model (2.1), together with the low-rank coefficient tensor at (2.2), is referred to as the Tucker tensor regression model.
For simplicity, we denote $q_{N+m}=p_{m}$, $s_{N+m}=r_{m}$ and $\boldsymbol{U}_{N+m}=\boldsymbol{V}_{m}$ for $1\leq m\leq M$. Note that the Tucker decomposition at (2.2) is not unique, since $\mbox{$\mathscr{A}$}=\mbox{$\mathscr{G}$}\times_{1}\boldsymbol{U}_{1}\times_{2}\cdots\times_{N+M}\boldsymbol{U}_{N+M}=\left(\mbox{$\mathscr{G}$}\times_{1}\boldsymbol{O}_{1}\times_{2}\cdots\times_{N+M}\boldsymbol{O}_{N+M}\right)\times_{1}\left(\boldsymbol{U}_{1}\boldsymbol{O}_{1}^{-1}\right)\times_{2}\cdots\times_{N+M}\left(\boldsymbol{U}_{N+M}\boldsymbol{O}_{N+M}^{-1}\right)$ for any invertible matrices $\boldsymbol{O}_{i}\in\mathbb{R}^{s_{i}\times s_{i}}$ with $1\leq i\leq N+M$. Without loss of generality, we restrict these factor matrices to be orthonormal, i.e., $\boldsymbol{U}_{j}\in\mathbb{O}^{q_{j}\times s_{j}}$ for each $1\leq j\leq N+M$; see, for example, the higher-order singular value decomposition (HOSVD) in kolda2009tensor. Consequently, the Tucker tensor regression model can be rewritten into
[…]
Let $\boldsymbol{U}=\boldsymbol{U}_{N}\otimes\dots\otimes\boldsymbol{U}_{1}\in\mathbb{O}^{Q\times s}$ and $\boldsymbol{V}=\boldsymbol{V}_{M}\otimes\cdots\otimes\boldsymbol{V}_{1}\in\mathbb{O}^{P\times r}$, where $Q\vcentcolon=\prod_{n=1}^{N}q_{n}$, $P\vcentcolon=\prod_{m=1}^{M}p_{m}$, $s\vcentcolon=\prod_{n=1}^{N}s_{n}$, and $r\vcentcolon=\prod_{m=1}^{M}r_{m}$. Note that $\boldsymbol{U}_{n}$’s and $\boldsymbol{V}_{m}$’s are still not unique, and hence $\boldsymbol{U}$ and $\boldsymbol{V}$. However, the corresponding column spaces $\text{col}(\boldsymbol{U})$ and $\text{col}(\boldsymbol{V})$, as well as their projection matrices $\boldsymbol{P}_{\boldsymbol{U}}=\boldsymbol{U}\boldsymbol{U}^{\top}$ and $\boldsymbol{P}_{\boldsymbol{V}}=\boldsymbol{V}\boldsymbol{V}^{\top}$, can be uniquely defined.
Note that model (2.3) involves dimension reduction on both the responses and predictors. We first consider responses, and project $\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$ onto $\text{col}(\boldsymbol{U})$ and its orthogonal complement, i.e., $\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})=\boldsymbol{P}_{\boldsymbol{U}}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})+\boldsymbol{P}_{\boldsymbol{U}}^{\perp}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$ with $\boldsymbol{P}_{\boldsymbol{U}}^{\perp}=\boldsymbol{I}_{Q}-\boldsymbol{U}\boldsymbol{U}^{\top}$. It then holds that
[…]
where all predictable information of $\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$ is contained in $\text{col}(\boldsymbol{U})$. Moreover,
$\mathrm{cov}\left\{\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t}),\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})|\boldsymbol{P}_{\boldsymbol{V}}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\right\}=\boldsymbol{0},$
indicating that $\text{col}(\boldsymbol{V})$ contains all useful information of predictors in terms of predicting responses. As a result, define $\boldsymbol{f}_{t}^{\mathrm{response}}=\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t}\times_{n=1}^{N}\boldsymbol{U}_{n}^{\top})=\boldsymbol{U}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})\in\mathbb{R}^{s}$ and $\boldsymbol{f}_{t}^{\mathrm{predictor}}=\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t}\times_{m=1}^{M}\boldsymbol{V}_{m}^{\top})=\boldsymbol{V}^{\top}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\in\mathbb{R}^{r}$ to be the response and predictor factors, respectively, and then the Tucker tensor regression can be well interpreted from a supervised factor modeling perspective (wang2022high; wang2024high; huang2025supervised).
The above supervised factor modeling interpretation also motivates us to introduce a new framework to construct a tensor regression model, especially for data with complex structures beyond the capability of Tucker decomposition, or for non-orthogonal loadings in the feature extraction process such as the CP decomposition (harshman1970foundations). Specifically, we first extract features or factors from responses and predictors by
$\boldsymbol{f}_{t}^{\mathrm{predictor}}=\mathcal{M}(\mbox{$\mathscr{X}$}_{t})\quad\text{and}\quad\boldsymbol{f}_{t}^{\mathrm{response}}=\mathcal{N}(\mbox{$\mathscr{Y}$}_{t}),$ (2.4)
respectively. Here, $\mathcal{M}(\cdot):\mathbb{R}^{p_{1}\times\cdots\times p_{M}}\to\mathbb{R}^{r}$ and $\mathcal{N}(\cdot):\mathbb{R}^{q_{1}\times\cdots\times q_{N}}\to\mathbb{R}^{s}$ are feature-extracting operators, with $r$ and $s$ being the number of extracted predictor and response features, respectively. A specific structure is then assumed to the coefficient tensor $\mathscr{A}$ of model (2.1) such that $\boldsymbol{f}_{t}^{\mathrm{response}}=\boldsymbol{\Theta}\boldsymbol{f}_{t}^{\mathrm{predictor}}+\boldsymbol{e}_{t}$, where $\boldsymbol{\Theta}\in\mathbb{R}^{s\times r}$ and $\boldsymbol{e}_{t}\in\mathbb{R}^{s}$ are the corresponding coefficient matrix and compressed error term, respectively. To this end, if operators $\mathcal{M}$ and $\mathcal{N}$ are linear, i.e., $\mathcal{M}(\mbox{$\mathscr{X}$}_{t})=\boldsymbol{\Lambda}_{x}^{\top}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$ and $\mathcal{N}(\mbox{$\mathscr{Y}$}_{t})=\boldsymbol{\Lambda}_{y}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$, with non-degenerating loading matrices $\boldsymbol{\Lambda}_{x}\in\mathbb{R}^{P\times r}$ and $\boldsymbol{\Lambda}_{y}\in\mathbb{R}^{Q\times s}$, we then can set the coefficient tensor to
$\mbox{$\mathscr{A}$}]_{N}=\boldsymbol{\Lambda}_{y}\left(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y}\right)^{-1}\boldsymbol{\Theta}\boldsymbol{\Lambda}_{x}^{\top}.$ (2.5)
Moreover, if $\boldsymbol{\Lambda}_{y}$ is orthonormal, then $\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y}=\boldsymbol{I}_{s}$ and $\mbox{$\mathscr{A}$}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}\boldsymbol{\Lambda}_{x}^{\top}$.
When Tucker decomposition is applied as in (2.2), the loading matrices $\boldsymbol{\Lambda}_{y}=\boldsymbol{U}$ and $\boldsymbol{\Lambda}_{x}=\boldsymbol{V}$ extract features simultaneously from all modes of the response and predictor, respectively. Such simultaneous compression may not be suitable for more complex data, say with hierarchical structures in the next subsection, and a new methodology is required.
2.2 Purely hierarchical structure
According to technological advances, the available multi-dimensional or tensor-valued data have become more and more complex, and it is very common to observe hierarchical relationships among some modes. In order to exploit their structural information, we attempt to design a regression model using the proposed framework at (2.4) and (2.5). This subsection starts from the simple case with purely hierarchical responses and predictors.
Figure 1: Three hierarchical feature extraction procedures for a second-order tensor $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}}$: hierarchical factor model (HFM), higher-order factor model (HOFM) and our model.
We first consider the feature extraction on predictors, and assume for simplicity that $M=2$, i.e., $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}}$, and the $p_{1}$ variables of the first mode are nested within each of the $p_{2}$ variables in its second mode. There are two commonly used methods in the literature of factor analysis to extract features from such hierarchical data $\mbox{$\mathscr{X}$}_{t}$: the hierarchical factor model (HFM) in schmid1957development and higher-order factor model (HOFM) in yung1999relationship. The HFM directly extracts features from each level of the hierarchy separately and jointly. In contrast, the HOFM performs feature extraction in a sequential manner along the hierarchy, where features extracted from a lower level are used to summarize those from the upper level. The loading matrices for these two methods are illustrated in Figure 1. Note that the HFM has more than $p_{2}$ features, while the HOFM has $O(rP)$ parameters with $P=p_{1}p_{2}$, making both of them infeasible in high-dimensional settings. Moreover, Tucker decomposition method extracts features simultaneously from all modes, i.e. $\mathcal{M}(\mbox{$\mathscr{X}$}_{t})=\operatorname*{vec}(\boldsymbol{V}_{1}^{\top}\mbox{$\mathscr{X}$}_{t}\boldsymbol{V}_{2})$, and it is not suitable to handle such hierarchical relationships.
Note that the HOFM employs different loadings for each level of the hierarchy. Such flexibility can maintain its performance, while the corresponding number of parameters is also enlarged dramatically for a larger dimension of $M$. This paper alternatively proposes a sequential feature extraction method with shared loadings within each level of the hierarchy. Specifically, we define the operator $\mathcal{M}(\mbox{$\mathscr{X}$}_{t};\{\boldsymbol{G}_{m}\}_{m=1}^{2})=\boldsymbol{G}_{2}^{\top}(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$, where $\boldsymbol{G}_{1}\in\mathbb{O}^{p_{1}\times r_{1}}$ and $\boldsymbol{G}_{2}\in\mathbb{O}^{p_{2}r_{1}\times r_{2}}$ are the component matrices for summarizing predictors, where $r_{1}$ and $r_{2}$ are the interim dimensions of features extracted from the corresponding mode. To interpret this construction, we liken the hierarchy of $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}}$ to a tree structure with $p_{2}$ branches, each branch containing $p_{1}$ leaves. For each branch, $\boldsymbol{G}_{1}^{\top}$ extracts $r_{1}$ features from its leaves to represent this branch. Overall, we have $r_{1}p_{2}$ features extracted from all leaves, given by $(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$. Then the subsequent $\boldsymbol{G}_{2}^{\top}\in\mathbb{R}^{r_{2}\times p_{2}r_{1}}$ combines the extracted features from all branches to form $r_{2}$ features for the entire tree; see Figure 1(c) for the illustration of this hierarchical feature extraction procedure. Therefore, the overall loading matrix connecting the $p_{1}p_{2}$ leaves to the final $r_{2}$ features for the tree is $\boldsymbol{G}_{2}^{\top}(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})$, and the column space of this loading can dwell in an arbitrary subspace of $\mathbb{R}^{r_{2}}$.
Remark 1.
Consider a special case with $p_{2}=2$, i.e., predictors $\mbox{$\mathscr{X}$}_{t}$ can be divided into two branches, $\boldsymbol{x}_{1:p_{1},1}$ and $\boldsymbol{x}_{1:p_{1},2}$. Suppose that the first branch has common features $\boldsymbol{G}^{*\top}\boldsymbol{x}_{1:p_{1},1}\in\mathbb{R}^{2}$ and branch-specific one $\boldsymbol{G}_{1}^{*\top}\boldsymbol{x}_{1:p_{1},1}\in\mathbb{R}$, while the second one has the common and branch-specific features of $\boldsymbol{G}^{*\top}\boldsymbol{x}_{1:p_{1},2}\in\mathbb{R}^{2}$ and $\boldsymbol{G}_{2}^{*\top}\boldsymbol{x}_{1:p_{1},2}\in\mathbb{R}^{2}$, respectively. Note that the HOFM uses loading matrices $\boldsymbol{G}_{1}^{(1)}\in\mathbb{R}^{p_{1}\times 3}$ and $\boldsymbol{G}_{1}^{(2)}\in\mathbb{R}^{p_{1}\times 4}$ to extract features from the first and second branches, respectively, while the proposed method employs a common loading matrix $\boldsymbol{G}_{1}\in\mathbb{R}^{p_{1}\times 5}$ to extract features from each branch. Shortly speaking, the HOFM concentrates more on branch-specific features, while our method focuses more on common features. In real applications, some weak branch-specific features may be ignored by the introduced method at the first layer of feature extraction, but they may also eventually be averaged out by the HOFM at later layers, i.e., our method does not miss globally important features.
For a general dimension of $M$, we assume that each mode of $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{M}}$ is nested within its subsequent mode, and the feature extraction can then be iterated until the last mode of $\mbox{$\mathscr{X}$}_{t}$. Specifically, for each $1\leq m\leq M$, we define the component matrix for summarizing predictors as $\boldsymbol{G}_{m}\in\mathbb{O}^{r_{m-1}p_{m}\times r_{m}}$, where $r_{0}=1$, $r_{m}$ is the interim dimension of features extracted from the corresponding mode, and $r_{M}=r$. As a result, the feature extraction on predictors at (2.4) can be formulated below,
$\boldsymbol{f}_{t}^{\mathrm{predictor}}=\mathcal{M}(\mbox{$\mathscr{X}$}_{t};\{\boldsymbol{G}_{m}\}_{m=1}^{M})=\underbrace{\boldsymbol{G}_{M}^{\top}(\boldsymbol{I}_{p_{M}}\otimes\boldsymbol{G}_{M-1}^{\top})\cdots(\boldsymbol{I}_{p_{2}\cdots p_{M}}\otimes\boldsymbol{G}_{1}^{\top})}_{\vcentcolon=\boldsymbol{\Lambda}_{x}^{\top}}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t}).$ (2.6)
Similarly, if the response tensor $\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{q_{1}\times q_{2}\times\cdots\times q_{N}}$ also has such purely hierarchical structure, the feature extraction on responses can be formulated into $\boldsymbol{f}_{t}^{\mathrm{response}}=\mathcal{N}(\mbox{$\mathscr{Y}$}_{t};\{\boldsymbol{H}_{n}\}_{n=1}^{N})=\boldsymbol{H}_{N}^{\top}(\boldsymbol{I}_{q_{N}}\otimes\boldsymbol{H}_{N-1}^{\top})\cdots(\boldsymbol{I}_{q_{2}\cdots q_{N}}\otimes\boldsymbol{H}_{1}^{\top})\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})\vcentcolon=\boldsymbol{\Lambda}_{y}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$, where $\boldsymbol{H}_{n}\in\mathbb{O}^{s_{n-1}q_{n}\times s_{n}}$ with $s_{0}=1$ and $s_{N}=s$ are the component matrices for $1\leq n\leq N$. Note that $\boldsymbol{\Lambda}_{y}$ is orthonormal and, for the tensor regression model at (2.1), its coefficient tensor $\mathscr{A}$ can then be defined as $\mbox{$\mathscr{A}$}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}\boldsymbol{\Lambda}_{x}^{\top}$ by adopting the method at (2.5), where $\boldsymbol{\Theta}\in\mathbb{R}^{s\times r}$ is a coefficient matrix.
Remark 2.
Consider a special case with $r=s$ and $\boldsymbol{\Theta}$ being diagonal, and then the above coefficient tensor $\mathscr{A}$ will admit a form of tensor train decomposition (oseledets2011tensor; zhou2022optimal; si2024efficient) if predictors are further arranged in a reverse hierarchical order. See Supplementary Material for the detailed illustration.
2.3 General hierarchical structure
This subsection discusses the case with partially hierarchical predictors and responses, and some of their modes may not be strictly nested within the others, possibly with unknown or even multiple hierarchical relationships among the modes.
We first consider the feature extraction on predictors. A natural way is first to assume a pseudo-full hierarchical structure to predictors, and the proposed method in the previous subsection can then be applied. The pseudo-structure should be consistent with the hierarchical structures intrinsic to the data, and the order of its modes is called an action order. For example, we consider the case with $M=3$, and mode 1 is nested within mode 2. The pseudo-structure can then be with an action order of $\{1,2,3\}$ or $\{3,1,2\}$.
In general, for the predictor tensor $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{M}}$, an action order can be denoted by $\alpha=(\alpha_{(1)},\alpha_{(2)},\ldots,\alpha_{(M)})$, which is a permutation of set $\{1,2,\ldots,M\}$. We then introduce an operator $\mathcal{T}(\cdot;\cdot)$ to permute $\mbox{$\mathscr{X}$}_{t}$ such that its modes follow the action order $\alpha$,
$\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha)_{i_{1}\ldots i_{M}}=(\mbox{$\mathscr{X}$}_{t})_{i_{\alpha_{(1)}}\dots i_{\alpha_{(M)}}}\hskip 8.53581pt\text{and}\hskip 8.53581pt\operatorname*{vec}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha))=\boldsymbol{T}_{x}(\alpha)\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t}),$
where $\boldsymbol{T}_{x}(\alpha)\in\mathbb{R}^{P\times P}$ is the permutation matrix. Note that $\boldsymbol{T}_{x}(\alpha)=\boldsymbol{I}_{P}$ when $\alpha=(1,2,\ldots,M)$. As a result, by applying (2.6), we then can extract features from predictors
$\boldsymbol{f}_{t}^{\mathrm{predictor}}=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha);\{\boldsymbol{G}_{m}\in\mathbb{O}^{r_{m-1}p_{\alpha_{(m)}}\times r_{m}}\}_{m=1}^{M})\in\mathbb{R}^{r}\hskip 8.53581pt\text{with}\hskip 8.53581pt(r_{0},r_{M})=(1,r).$ (2.7)
Denote by $\mathcal{P}_{x}(M)$ the set of all possible action orders for predictors $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{M}}$. Assume that $\mbox{$\mathscr{X}$}_{t}$ has $m$ hierarchical structures, which occupy $M^{\prime}$ modes in total, and the set $\mathcal{P}_{x}(M)$ can then be verified to have $(M-M^{\prime}+m)!$ possible action orders.
Proposition 1.
For any $\mbox{$\mathscr{X}$}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{M}}$, an action order $\alpha=(\alpha_{(1)},\alpha_{(2)},\ldots,\alpha_{(M)})$ is used for the permutation, and factor matrices $\{\boldsymbol{G}_{m}\in\mathbb{O}^{r_{m-1}p_{\alpha_{(m)}}\times r_{m}}\}_{m=1}^{M}$ with $(r_{0},r_{M})=(1,r)$ are applied to extract features as in (2.7). It then holds that, for a different action order $\alpha^{\prime}=(\alpha_{(1)}^{\prime},\alpha_{(2)}^{\prime},\ldots,\alpha_{(M)}^{\prime})$, there exist $\{\boldsymbol{G}_{m}^{\prime}\in\mathbb{O}^{r^{\prime}_{m-1}p_{\alpha^{\prime}_{(m)}}\times r^{\prime}_{m}}\}_{m=1}^{M}$ with $(r_{0}^{\prime},r_{M}^{\prime})=(1,r)$ such that $\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$};\alpha);\{\boldsymbol{G}_{m}\}_{m=1}^{M})=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$};\alpha^{\prime});\{\boldsymbol{G}_{m}^{\prime}\}_{m=1}^{M})$.
The above proposition proves a nice equivalence of all elements inside $\mathcal{P}_{x}(M)$, however, it is a theoretical result only. In real applications, each action order may have a totally different number of parameters, and a wrongly selected order may end up with a huge number of parameters. As a result, we may choose the one with the smallest size,
$\alpha^{*}=\operatorname*{arg\,min}_{\alpha\in\mathcal{P}_{x}(M)}d_{x}(\alpha)\hskip 8.53581pt\text{with}\hskip 8.53581ptd_{x}(\alpha)=\sum_{m=1}^{M}r_{m-1}r_{m}p_{m}.$ (2.8)
On the other hand, for a certain predictor tensor $\mbox{$\mathscr{X}$}_{t}$, more than one action orders may be more practical, while the summarized features $\boldsymbol{f}_{t}^{\mathrm{predictor}}$ have a fixed dimension of $r$ regardless which action orders are used. This motivates us to choose the best set of action orders,
$\mathcal{P}_{x}(M,K_{x})\vcentcolon=\{\alpha_{1}^{*},\ldots,\alpha_{K_{x}}^{*}\}=\operatorname*{arg\,min}_{1\leq K\leq|\mathcal{P}_{x}(M)|,\alpha_{k}\in\mathcal{P}_{x}(M)}\sum_{k=1}^{K}d_{x}(\alpha_{k})\hskip 8.53581pt\text{subject to}\hskip 8.53581pt\sum_{k=1}^{K}r_{M}(\alpha_{k})=r,$ (2.9)
where $|\mathcal{P}_{x}(M)|$ is the cardinality of $\mathcal{P}_{x}(M)$, and $r_{M}(\alpha_{k})$ is the rank of $r_{M}$ with respect to $\alpha_{k}$. Note that the number of parameters for $\mathcal{P}_{x}(M,K_{x})$ is smaller than that for $\alpha^{*}$ from (2.8).
For the selected $\alpha^{*}$ from (2.8), we can extract features from predictors as in (2.7). This paper focuses on $\mathcal{P}_{x}(M,K_{x})$, the selected set of action orders. For each $1\leq k\leq K_{x}$, denote $\alpha_{k}^{*}=(\alpha_{(1)}^{k},\alpha_{(2)}^{k},\ldots,\alpha_{(M)}^{k})$ with ranks $(r_{0,k},r_{1,k},\ldots,r_{M,k})$, and let $\boldsymbol{f}_{k,t}^{\mathrm{predictor}}=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha_{k}^{*});\{\boldsymbol{G}_{m,k}\in\mathbb{O}^{r_{m-1,k}p_{\alpha^{k}_{(m)}}\times r_{m,k}}\}_{m=1}^{M})\in\mathbb{R}^{r_{M,k}}$. As a result, the extracted features from predictors can be denoted by $\boldsymbol{f}_{t}^{\mathrm{predictor}}=\left(\boldsymbol{f}_{1,t}^{\mathrm{predictor}\top},\ldots,\boldsymbol{f}_{K_{x},t}^{\mathrm{predictor}\top}\right)^{\top}=\boldsymbol{\Lambda}_{x}^{\top}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\in\mathbb{R}^{r}$, where $r=\sum_{k=1}^{K_{x}}r_{M,k}$, and the loading matrix $\boldsymbol{\Lambda}_{x}$ has the form of
$\boldsymbol{\Lambda}_{x}=\begin{bmatrix}\boldsymbol{G}_{M,1}^{\top}(\boldsymbol{I}_{p_{\alpha^{1}_{(M)}}}\otimes\boldsymbol{G}_{M-1,1}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{1}_{(2)}}\cdots p_{\alpha^{1}_{(M)}}}\otimes\boldsymbol{G}_{1,1}^{\top})\boldsymbol{T}_{x}(\alpha_{1})\\ \boldsymbol{G}_{M,2}^{\top}(\boldsymbol{I}_{p_{\alpha^{2}_{(M)}}}\otimes\boldsymbol{G}_{M-1,2}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{2}_{(2)}}\cdots p_{\alpha^{2}_{(M)}}}\otimes\boldsymbol{G}_{1,2}^{\top})\boldsymbol{T}_{x}(\alpha_{2})\\ \cdots\\ \boldsymbol{G}_{M,K_{x}}^{\top}(\boldsymbol{I}_{p_{\alpha^{K_{x}}_{(M)}}}\otimes\boldsymbol{G}_{M-1,K_{x}}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{K_{x}}_{(2)}}\cdots p_{\alpha^{K_{x}}_{(M)}}}\otimes\boldsymbol{G}_{1,K_{x}}^{\top})\boldsymbol{T}_{x}(\alpha_{K_{x}})\end{bmatrix}^{\top}$. (2.10)
The above matrix is no longer orthogonal in general. The philosophy of the above feature extraction is analogous to that of CP decomposition, which concatenates multiple separately extracted features. Note that, comparing with Tucker decomposition, CP decomposition requests much less parameters by sacrificing orthogonality of loading matrices.
The feature extraction from responses $\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{q_{1}\times q_{2}\times\cdots\times q_{N}}$ can be conducted in a similar way. Denote the corresponding action order by $\beta=(\beta_{(1)},\beta_{(2)},\ldots,\beta_{(N)})$, which is a permutation of set $\{1,2,\ldots,N\}$, and let $\mathcal{P}_{y}(N)$ be the set of all possible action orders. By a method similar to (2.8) and (2.9), we can define the best action order $\beta^{*}$ and the best set of action orders $\mathcal{P}_{y}(N,K_{y})\vcentcolon=\{\beta_{1}^{*},\ldots,\beta_{K_{y}}^{*}\}=\operatorname*{arg\,min}_{1\leq K\leq|\mathcal{P}_{y}(N)|,\beta_{k}\in\mathcal{P}_{y}(N)}\sum_{k=1}^{K}d_{y}(\beta_{k})$ with $d_{y}(\beta)=\sum_{n=1}^{N}s_{n-1}s_{n}q_{n}$, respectively. Moreover, for each $1\leq k\leq K_{y}$, let $\beta_{k}^{*}=(\beta_{(1)}^{k},\beta_{(2)}^{k},\ldots,\beta_{(N)}^{k})$ with ranks $(s_{0,k},s_{1,k},\ldots,s_{N,k})$, component matrices be $\boldsymbol{H}_{n,{k}}\in\mathbb{O}^{s_{n-1,{k}}q_{\beta^{k}_{(n)}}\times s_{n,{k}}}$ with $1\leq n\leq N$, and the corresponding extracted features be $\boldsymbol{f}_{{k},t}^{\mathrm{response}}=\mathcal{N}(\mathcal{T}(\mbox{$\mathscr{Y}$}_{t};\beta^{*}_{k});\{\boldsymbol{H}_{n,{k}}\}_{n=1}^{N})$. As a result, the total extracted features from responses can be represented as $\boldsymbol{f}_{t}^{\mathrm{response}}=\left(\boldsymbol{f}_{1,t}^{\mathrm{response}\top},\ldots,\boldsymbol{f}_{K_{y},t}^{\mathrm{response}\top}\right)^{\top}\vcentcolon=\boldsymbol{\Lambda}_{y}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})\in\mathbb{R}^{s}$, where $s=\sum_{k=1}^{K_{y}}s_{N,k}$, and the loading matrix $\boldsymbol{\Lambda}_{y}$ can be defined similar to (2.10) with the permutation matrix $\boldsymbol{T}_{y}(\beta)\in\mathbb{R}^{Q\times Q}$.
From (2.5), the coefficient tensor of model (2.1) can be designed to have the form of
$[\mbox{$\mathscr{A}$}_{\mathrm{LR}}]_{N}=\boldsymbol{\Lambda}_{y}\left(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y}\right)^{-1}\boldsymbol{\Theta}_{\mathrm{LR}}\boldsymbol{\Lambda}_{x}^{\top}$ (2.11)
with $\boldsymbol{\Theta}_{\mathrm{LR}}\in\mathbb{R}^{s\times r}$ being the coefficient matrix, leading to an efficient and interpretable tensor regression model for data with hierarchical structures.
2.4 Autoregressive model for hierarchical time series
Consider an autoregressive model for hierarchical time series $\{\mbox{$\mathscr{Y}$}_{t}\}$ with $\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{q_{1}\times q_{2}\times\cdots\times q_{N}}$,
[…]
Assumption 1.
The determinant of the matrix polynomial $\mbox{$\mathscr{A}$}(z)=\boldsymbol{I}_{Q}-\boldsymbol{A}_{1}z-\cdots-\boldsymbol{A}_{L}z^{L}$ is not equal to zero for all $z\in\mathbb{C}$ and $|z|<1$.
We first consider the case with $L=1$, and the coefficient tensor $\mbox{$\mathscr{A}$}_{1}$ can be directly assumed to have the decomposition in (2.11) with the same definitions of $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$, where $M=N$, $p_{n}=q_{n}$ for $1\leq n\leq N$, and $\mbox{$\mathscr{X}$}_{t}=\mbox{$\mathscr{Y}$}_{t-1}$. As a result, model (2.12) can then be rewritten into an interpretable form:
[…]
Note that the high-order tensor $\mbox{$\mathscr{Y}$}_{t}$ is summarized into features $\boldsymbol{f}_{t}^{\mathrm{response}}\in\mathbb{R}^{s}$, while the information from $\mbox{$\mathscr{Y}$}_{t-1}$ is compressed into $\boldsymbol{f}^{\mathrm{predictor}}_{t}\in\mathbb{R}^{r}$. The coefficient matrix $\boldsymbol{\Theta}_{1}\in\mathbb{R}^{s\times r}$ characterizes the linear relationship between features $\boldsymbol{f}^{\mathrm{response}}_{t}$ and $\boldsymbol{f}^{\mathrm{predictor}}_{t}$.
For a general order $L$, we apply (2.5) to each coefficient tensor $\mbox{$\mathscr{A}$}_{l}$, resulting in:
$\mbox{$\mathscr{A}$}_{l}]_{N}=\boldsymbol{\Lambda}_{y}\left(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y}\right)^{-1}\boldsymbol{\Theta}_{l}\boldsymbol{\Lambda}_{x}^{\top}\quad\text{for}\quad 1\leq l\leq L.$ (2.13)
[…]
3 High-dimensional estimation
3.1 Ordinary least squares approach
This paper considers the ordinary least squares (OLS) method to estimate the proposed tensor autoregressive model at (2.12) and (2.13) with hierarchical time series.
Suppose that both sets of action orders $\mathcal{P}_{x}(M,K_{x})=\{\alpha_{1},\ldots,\alpha_{K_{x}}\}$ and $\mathcal{P}_{y}(N,K_{y})=\{\beta_{1},\ldots,\beta_{K_{y}}\}$ are known and, for all $1\leq k\leq K_{x}$ and $1\leq k^{\prime}\leq K_{y}$, the corresponding ranks $\{r_{m,k}\}_{m=1}^{M}$ and $\{s_{n,k^{\prime}}\}_{n=1}^{N}$ are also given. Let $\boldsymbol{p}=(p_{1},\ldots,p_{M})$, $\boldsymbol{q}=(q_{1},\ldots,q_{N})$, $\boldsymbol{r}_{k}=(r_{1,k},\ldots,r_{M,k})$ and $\boldsymbol{s}_{k^{\prime}}=(s_{1,k^{\prime}},\ldots,s_{N,k^{\prime}})$, where $1\leq k\leq K_{x}$ and $1\leq k^{\prime}\leq K_{y}$. Note that $\sum_{k=1}^{K_{x}}r_{M,k}=r$, $\sum_{k=1}^{K_{y}}s_{N,k}=s$, $M=N$ and $\boldsymbol{p}=\boldsymbol{q}$. We then denote the parameter spaces of loading matrices from predictor and response features respectively by
[…]
For autoregressive models at (2.12) and (2.13), the coefficient tensor can be denoted by $\mbox{$\mathscr{A}$}_{\mathrm{AR}}\in\mathbb{R}^{q_{1}\times\cdots\times q_{N}\times q_{1}\times\cdots\times q_{N}\times L}$ with $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=([\mbox{$\mathscr{A}$}_{1}]_{N},[\mbox{$\mathscr{A}$}_{2}]_{N},\ldots,[\mbox{$\mathscr{A}$}_{L}]_{N})$ being the mode-$N$ sequential matricization. Let $\boldsymbol{\Theta}_{\mathrm{AR}}=[\boldsymbol{\Theta}_{1},\ldots,\boldsymbol{\Theta}_{L}]\in\mathbb{R}^{s\times rL}$, and it holds that $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=\boldsymbol{\Lambda}_{y}(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y})^{-1}\boldsymbol{\Theta}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$. Moreover, by reparameterizing $(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y})^{-1}\boldsymbol{\Theta}_{\mathrm{AR}}$ into $\boldsymbol{\Theta}_{\mathrm{AR}}$, we have $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$, and then the parameter space can be defined as
$\begin{split}\mathcal{S}_{\mathrm{AR}}=\{\mbox{$\mathscr{A}$}\in&\mathbb{R}^{q_{1}\times\cdots\times q_{N}\times p_{1}\times\cdots\times p_{M}\times L}:\penalty 10000\ [\mbox{$\mathscr{A}$}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top}),\penalty 10000\ \boldsymbol{\Theta}\in\mathbb{R}^{s\times rL},\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \boldsymbol{\Lambda}_{y}\in\mathcal{R}_{y}(\boldsymbol{q},\{\boldsymbol{s}_{k}\}_{k=1}^{K_{y}},\mathcal{P}_{y}(N,K_{y})),\penalty 10000\ \boldsymbol{\Lambda}_{x}\in\mathcal{R}_{x}(\boldsymbol{p},\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}},\mathcal{P}_{x}(M,K_{x}))\}.\end{split}$
For an observed time series $\{\mbox{$\mathscr{Y}$}_{1-L},\ldots,\mbox{$\mathscr{Y}$}_{0},\mbox{$\mathscr{Y}$}_{1},\ldots,\mbox{$\mathscr{Y}$}_{T}\}$ generated by the model at (2.12) and (2.13), the predictor tensor $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{q_{1}\times\cdots\times q_{N}\times L}$ can be formed by stacking $\mbox{$\mathscr{Y}$}_{t-1},\dots,\mbox{$\mathscr{Y}$}_{t-L}$ together, and the OLS estimator is hence defined as
$\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}=\operatorname*{arg\,min}_{\mbox{$\mathscr{A}$}\in\mathcal{S}_{\mathrm{AR}}}\mathcal{L}_{T}(\mbox{$\mathscr{A}$})\quad\text{with}\quad\mathcal{L}_{T}(\mbox{$\mathscr{A}$})=\frac{1}{T}\sum_{t=1}^{T}\left\lVert\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})-[\mbox{$\mathscr{A}$}]_{N}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\right\rVert_{2}^{2}.$
Moreover, for the tensor regression model at (2.1) and (2.11), by letting $L=1$, we can similarly define its parameter space and OLS estimator, denoted by $\mathcal{S}_{\mathrm{LR}}$ and $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$, respectively.
Denote by $\mbox{$\mathscr{A}$}^{*}$ the true coefficient tensor of the tensor regression or autoregressive model in the previous section, and it depends on the best set of action orders, which is unknown in real applications. In fact, $\mathcal{P}_{x}(M,K_{x})$ and $K_{x}$ in this section are only the running set and numbers of action orders for predictors, respectively, and the situation is the same for $\mathcal{P}_{y}(N,K_{y})$ and $K_{y}$. Fortunately, from Proposition 1, we can project the true coefficient tensor $\mbox{$\mathscr{A}$}^{*}$ into the running sets $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$, with running numbers $K_{x}$ and $K_{y}$. Such projection has no error, while the number of parameters will be enlarged. Without loss of generality, we will assume $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$ to be the best sets in establishing theoretical properties, and Section 4 will discuss the selection of these hyperparameters.
3.2 Non-asymptotic properties for tensor regression
[…]
We begin with the tensor-on-tensor regression setting, where $\{(\mbox{$\mathscr{Y}$}_{t},\mbox{$\mathscr{X}$}_{t})\}_{t=1}^{T}$ are assumed to be independent and identically distributed (i.i.d.) observations. The following assumptions are imposed.
Assumption 2.
Predictors $\{\boldsymbol{x}_{t}\}$ are i.i.d. with $\mathbb{E}(\boldsymbol{x}_{t})=\boldsymbol{0}$, $\mathrm{var}(\boldsymbol{x}_{t})=\boldsymbol{\Sigma}_{x}$, and follow a $\sigma^{2}$-sub-Gaussian distribution. There exist constants $0<c_{x}<C_{x}<\infty$ such that $c_{x}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{x})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{x})\leq C_{x}$, where the two quantities $c_{x}$ and $C_{x}$ depend on the dimensions of $p_{j}$’s, and they may shrink to zero or diverge to infinity as the dimensions increase.
Assumption 3.
Error terms $\{\boldsymbol{e}_{t}\}$ are i.i.d. and, conditional on $\boldsymbol{x}_{t}$, $\boldsymbol{e}_{t}$ has mean zero and follows $\kappa^{2}$-sub-Gaussian distribution.
Assumption 4.
The matrix $\boldsymbol{\Theta}_{\mathrm{LR}}$ has its operator norm upper bounded by a constant $g_{1}>0$.
Assumption 2 guarantees the well-conditioning of the covariance matrix of the predictors. Assumption 3, on the other hand, controls the tail behavior of the noise by assuming sub-Gaussianity, a standard requirement in high-dimensional statistics that facilitates concentration inequalities. Both assumptions are widely adopted in the literature on high-dimensional regression and low-rank estimation; see, for example, wainwright2019high; si2024efficient; cai2025efficient. Assumption 4 imposes an upper bound on the operator norm of the coefficient matrix $\boldsymbol{\Theta}_{\mathrm{LR}}$. This is a mild technical condition, as the dimensions $s_{\mathrm{tot}}$ and $r_{\mathrm{tot}}$ of $\boldsymbol{\Theta}_{\mathrm{LR}}$ are typically small by carefully choosing the action orders.
The following theorem provides a non-asymptotic upper bound on the estimation error of estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$.
Theorem 1.
Suppose that Assumptions 2 - 4 hold, and the true coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{LR}}^{*}\in\mathcal{S}_{\mathrm{LR}}$. If the sample size $T\gtrsim(\sigma^{4}/c_{x}^{2})d_{\mathrm{LR}}\log((\sigma^{4}/c_{x}^{2})d_{\mathrm{LR}})$, then there exists $c_{1}>0$ such that
$\|\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{LR}}-\mbox{$\mathscr{A}$}^{*}_{\mathrm{LR}}\|_{\mathrm{F}}\lesssim\frac{\kappa\sigma}{c_{x}}\sqrt{\frac{d_{\mathrm{LR}}\log T}{T}}$
with probability at least $1-2\exp\left\{-c_{1}d_{\mathrm{LR}}\log T\right\}$, where $d_{\mathrm{LR}}=\sum_{k=1}^{K_{x}}d_{x}(\alpha_{k})+\sum_{k=1}^{K_{y}}d_{y}(\beta_{k})+sr$.
We define $d_{\mathrm{LR}}$ as the model complexity of the low-rank regression estimator, which can be rewritten explicitly as $d_{\mathrm{LR}}=K_{x}r_{\max}^{2}\sum_{m=1}^{M}p_{m}+K_{y}s_{\max}^{2}\sum_{n=1}^{N}q_{n}+K_{x}K_{y}r_{\max}s_{\max}$, where $r_{\max}\vcentcolon=\max_{m,k}r_{m,k}$ and $s_{\max}\vcentcolon=\max_{n,k^{\prime}}s_{n,k^{\prime}}$. From Theorem 1, our proposed low-rank regression estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$, up to a logarithmic factor, achieves an estimation error bound of order $O_{p}(\sqrt{d_{\mathrm{LR}}/T})$ when the sample size $T\gtrsim d_{\mathrm{LR}}/\text{SNR}^{2}$, where $\text{SNR}=c_{x}/\sigma^{2}$ is the predictor signal-to-noise ratio. This rate aligns with the general optimal rate for low-rank regression established in han2022optimal. The model complexity $d_{\mathrm{LR}}$ grows linearly with the dimensions of each mode $\{p_{m}\}$ and $\{q_{n}\}$, and quadratically with the interim ranks $\{r_{m,k}\}$ and $\{s_{n,k^{\prime}}\}$. The logarithmic factor $\log T$, negligible in most common cases, arises from the non-orthogonality of the loadings. In contrast, if we assume the coefficient tensor $\mathscr{A}$ follows a Tucker decomposition with Tucker ranks $(s_{1},s_{2},\ldots,s_{N},r_{1},r_{2},\ldots,r_{M})$, the corresponding complexity would be $d_{\mathrm{Tucker}}=\prod_{n=1}^{N}s_{n}\times\prod_{m=1}^{M}r_{m}+\sum_{n=1}^{N}q_{n}s_{n}+\sum_{m=1}^{M}p_{m}r_{m}$, which grows exponentially with the tensor orders $N$ and $M$ (han2022optimal; wang2022high).
3.3 Non-asymptotic properties for tensor autoregression
We now turn to the tensor autoregressive model specified in (2.12). This setting introduces temporal dependence, and the assumptions are adapted accordingly.
Assumption 5.
Let error terms $\boldsymbol{e}_{t}=\boldsymbol{\Sigma}_{e}^{1/2}\boldsymbol{\xi}_{t}$. Random vectors $\{\boldsymbol{\xi}_{t}\}$ are i.i.d. with mean zero and covariance $\boldsymbol{I}_{Q}$, the entries $(\boldsymbol{\xi}_{tj})_{1\leq j\leq Q}$ of $\boldsymbol{\xi}_{t}$ are mutually independent and $\kappa^{2}$-sub-Gaussian distributed, and there exist constants $0<c_{e}<C_{e}<\infty$ such that $c_{e}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{e})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{e})\leq C_{e}$.
Assumption 6.
The matrix $\boldsymbol{\Theta}_{\mathrm{AR}}$ has its operator norm upper bounded by a constant $g_{2}>0$.
The sub-Gaussianity assumption in Assumption 5 is commonly used in the analysis of high-dimensional time series models. It provides control over the tail behavior of the error terms and facilitates the use of concentration inequalities. This assumption is widely adopted in recent literature, including zheng2020finite; wang2024high; huang2025supervised. Assumption 6 is essential for the consistency of the estimator in the autoregressive setting, which is relatively mild and practically reasonable. In our setting, $\boldsymbol{\Lambda}_{y}$ and $\boldsymbol{\Lambda}_{x}$ may exhibit multi-collinearity, and this assumption is weaker than enforcing a lower bound on the smallest singular value of $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$ for model well-behavedness.
Following basu2015regularized, we define the following quantities,
$\mu_{\mathrm{min}}(\mbox{$\mathscr{A}$})=\min_{|z|=1}\lambda_{\mathrm{min}}(\bar{\mbox{$\mathscr{A}$}}(z)\mbox{$\mathscr{A}$}(z))\quad\text{and}\quad\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$})=\max_{|z|=1}\lambda_{\mathrm{max}}(\bar{\mbox{$\mathscr{A}$}}(z)\mbox{$\mathscr{A}$}(z)),$
where $\bar{\mbox{$\mathscr{A}$}}(z)$ denotes the conjugate transpose of $\mbox{$\mathscr{A}$}(z)$.
Furthermore, the AR($L$) model in (2.12) can be alternatively represented as a $QL$-dimensional AR(1) model with its transition matrix $\boldsymbol{B}$ defined in (C.3) in the Supplementary Material. We similarly define the matrix polynomial associated with the lifted AR(1) representation as $\boldsymbol{B}(z)=\boldsymbol{I}_{QL}-\boldsymbol{B}z$. We then define
$\mu_{\mathrm{min}}(\boldsymbol{B})=\min_{|z|=1}\lambda_{\mathrm{min}}(\bar{\boldsymbol{B}}(z)\boldsymbol{B}(z))\quad\text{and}\quad\mu_{\mathrm{max}}(\boldsymbol{B})=\max_{|z|=1}\lambda_{\mathrm{max}}(\bar{\boldsymbol{B}}(z)\boldsymbol{B}(z)),$
which quantify the smallest and largest eigenvalues the Hermitian matrix $\bar{\boldsymbol{B}}(z)\boldsymbol{B}(z)$ can take on the unit circle. Note that $\mu_{\mathrm{min}}(\boldsymbol{B})$, $\mu_{\mathrm{max}}(\boldsymbol{B})$ are not necessarily the same as $\mu_{\mathrm{min}}(\mbox{$\mathscr{A}$})$, $\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$})$ (basu2015regularized), and we can show $\mu_{\mathrm{min}}(\boldsymbol{B}^{\ast})\leq\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$.
We then state the error upper bound for the estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}$ below.
Theorem 2.
Suppose that Assumptions 1, 5, and 6 hold, and the true coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}\in\mathcal{S}_{\mathrm{AR}}$. If the sample size $T\gtrsim(\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2})d_{\mathrm{AR}}\log((\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2})d_{\mathrm{AR}})$, then
$\|{\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}-\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}}\|_{\mathrm{F}}\lesssim\frac{\kappa^{2}\kappa_{U,B}^{\prime}}{\kappa_{L,A}}\sqrt{\frac{d_{\mathrm{AR}}\log T}{T}}$
with probability at least $1-2\exp\left\{-c_{1}d_{\mathrm{AR}}\log T\right\}$, where $d_{\mathrm{AR}}=\sum_{k=1}^{K_{x}}d_{x}(\alpha_{k})+\sum_{k=1}^{K_{y}}d_{y}(\beta_{k})+Lsr$, $\boldsymbol{B}^{\ast}$ is the true value of $\boldsymbol{B}$ constructed from $\mbox{$\mathscr{A}$}^{\ast}$, $\kappa_{U,B}=C_{e}/\mu_{\mathrm{min}}(\boldsymbol{B}^{\ast})$, $\kappa_{L,A}=c_{e}/\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$, $\kappa^{\prime}_{U,B}=C_{e}/\mu^{1/2}_{\mathrm{min}}(\boldsymbol{B}^{\ast})$, and $c_{1}>0$ is a constant.
We define $d_{\mathrm{AR}}$ as the model complexity of the low-rank autoregressive estimator, which can be explicitly expressed as $d_{\mathrm{AR}}=(K_{x}r_{\max}^{2}+K_{y}s_{\max}^{2})\sum_{n=1}^{N}q_{n}+LK_{x}K_{y}r_{\max}s_{\max}$. When the order $N$, the number of action orders $K_{x}$, $K_{y}$, and the largest ranks $s_{\max}$ and $r_{\max}$ are held fixed, the key parameters $c_{e}$, $C_{e}$, $\kappa^{2}$, $\mu_{\mathrm{min}}(\boldsymbol{B}^{\ast})$, and $\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$ are all bounded away from zero and infinity. Consequently, the estimation error of the low-rank tensor autoregressive estimator satisfies $\|{\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}-\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}}\|_{\mathrm{F}}=O_{p}(\sqrt{d_{\mathrm{AR}}/T})$ up to a logarithmic term, where $d_{\mathrm{AR}}$ is the effective model complexity of the autoregressive estimator.
4 Algorithms and hyperparameter selection
4.1 Algorithms
This subsection introduces an algorithm for computing the low-rank estimator defined in (3.1) under the general setting. To preserve the sequential low-rank structure of $\mbox{$\mathscr{A}$}]_{N}$, we adopt an alternating least squares (ALS) approach, which iteratively updates each component $\boldsymbol{H}_{n,k^{\prime}}$ and $\boldsymbol{G}_{m,k}$ while keeping all other components fixed.
We introduce some shorthand notations for clarity. Let $\boldsymbol{y}_{t}=\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$ and $\boldsymbol{x}_{t}=\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})=(\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t-1})^{\top},\dots,\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t-L})^{\top})^{\top}$, and our model can be rewritten in the vectorized form as:
$\boldsymbol{y}_{t}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})\boldsymbol{x}_{t}+\boldsymbol{e}_{t},$ (4.1)
[…]
We sequentially update each orthogonal matrix in each row block of $\boldsymbol{\Lambda}_{x}^{\top}$ (i.e., update $\boldsymbol{G}_{1,1}$ up to $\boldsymbol{G}_{M,1}$ for the first row block), and then switch to the next row block for another sequence of updates, and so on. Specifically, we convert the update of each component into a regression problem relegated to Supplementary Material, similar to coordinate descent. The update for $\boldsymbol{\Lambda}_{y}^{\top}$ follows a similar procedure. After updating all components in $\boldsymbol{\Lambda}_{x}^{\top}$ and $\boldsymbol{\Lambda}_{y}^{\top}$, we orthogonalize each row block of the two loading matrices and update the core tensor $\boldsymbol{\Theta}$ via an SSVD algorithm detailed in Supplementary Material. This process is repeated until convergence. For convenience, we arrange these matrices into $\boldsymbol{L}_{1},\boldsymbol{L}_{2},\ldots,\boldsymbol{L}_{K_{x}M+K_{y}N}$, where $\boldsymbol{L}_{i}$ is the $i$-th orthogonal matrix in the update procedure, and we re-define the loss function with regard to $\boldsymbol{L}_{i}$’s and $\boldsymbol{\Theta}$ as follows:
$\mathcal{L}_{T}\big(\boldsymbol{L}_{1},\dots,\boldsymbol{L}_{K_{x}M+K_{y}N},\boldsymbol{\Theta}\big)=\frac{1}{T}\sum_{t=1}^{T}\|\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})-\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})\operatorname*{vec}{(\mbox{$\mathscr{X}$}_{t})}\|^{2},$
$\text{s.t., }\boldsymbol{G}_{m,k}=\boldsymbol{L}_{(k-1)M+m}\quad\text{and}\quad\boldsymbol{H}_{n,k^{\prime}}=\boldsymbol{L}_{K_{x}M+(k^{\prime}-1)N+n}$ (4.2)
for all $m,k,n,k^{\prime}$, and $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$ are defined in (2.10), respectively. The overall algorithm is summarized below:
Algorithm 1 ALS for low-rank estimation
Input: Data $\{(\boldsymbol{y}_{t},\boldsymbol{x}_{t})\}_{t=1}^{T}$, initial coefficient $\mbox{$\mathscr{A}$}^{(0)}]_{N}=\boldsymbol{\Lambda}_{y}^{(0)}\boldsymbol{\Theta}^{(0)}\boldsymbol{\Lambda}_{x}^{(0)\top}$, and components $\{\{\boldsymbol{H}^{(0)}_{n,k^{\prime}}\}_{n=1}^{N}\}_{k^{\prime}=1}^{K_{y}}$ and $\{\{\boldsymbol{G}^{(0)}_{m,k}\}_{m=1}^{M}\}_{k=1}^{K_{x}}$ arranged into $\{\boldsymbol{L}_{i}^{(0)}\}_{i=1}^{K_{x}M+K_{y}N}$.
repeat
  for $i=1,2,\dots,K_{x}M+K_{y}N$ do
   $\boldsymbol{L}^{(u+0.5)}_{i}=\operatorname*{arg\,min}_{\boldsymbol{L}_{i}}\mathcal{L}_{T}\big(\boldsymbol{L}_{1}^{(u+0.5)},\dots,\boldsymbol{L}_{i-1}^{(u+0.5)},\boldsymbol{L}_{i},\boldsymbol{L}_{i+1}^{(u)},\dots,\boldsymbol{L}_{K_{x}M+K_{y}N}^{(u)},\boldsymbol{\Theta}^{(u)}\big)$
  end for
  $\{\boldsymbol{L}_{i}^{(u+1)}\}_{i=1}^{K_{x}M+K_{y}N},\boldsymbol{\Theta}^{(u)}=\text{SSVD}(\{\boldsymbol{L}_{i}^{(u+0.5)}\}_{i=1}^{K_{x}M+K_{y}N},\boldsymbol{\Theta}^{(u)})$
  $\boldsymbol{\Theta}^{(u+1)}=\operatorname*{arg\,min}_{\boldsymbol{\Theta}}\mathcal{L}_{T}\big(\boldsymbol{L}_{1}^{(u+1)},\dots,\boldsymbol{L}_{K_{x}M+K_{y}N}^{(u+1)},\boldsymbol{\Theta}\big)$
until convergence or $u=u_{\max}$, where $u_{\max}$ is the maximum iteration
Return $\mbox{$\mathscr{A}$}^{(u)}]_{N}=\boldsymbol{\Lambda}_{y}^{(u)}\boldsymbol{\Theta}^{(u)}\boldsymbol{\Lambda}_{x}^{(u)\top}$ via (4.2), (2.10).
4.2 Hyperparameter selection
Some hyperparameters involved in Algorithm 1 are typically unknown in most real-world applications, and their selection not only impacts the model’s performance, but also reveals the underlying structural information of the data. This subsection introduces a boosting-based procedure for determining the number of action orders $K_{y}$ and $K_{x}$, the corresponding action sets $\mathcal{P}_{y}(N,K_{y})$ and $\mathcal{P}_{x}(M,K_{x})$, and the associated rank sequences $\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}}$ and $\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}}$.
To guide the selection of these hyperparameters, we adopt an information criterion. Let $\mbox{$\mathscr{\widehat{A}}$}(R)$ denote the estimator from (3.1), where the configuration $R$ collects all hyperparameters: $R=(K_{y},K_{x},\mathcal{P}_{y}(N,K_{y}),\mathcal{P}_{x}(M,K_{x}),\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}},\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}})$. The Bayesian information criterion (BIC) is then constructed below,
$\mathrm{BIC}(R)=T\log\left\{\mathcal{L}_{T}(\mbox{$\mathscr{\widehat{A}}$}(R))\right\}+\phi d(R)\log(T),$ (4.3)
where $\phi$ is a tuning parameter, and $d(R)=\sum_{k^{\prime}=1}^{K_{y}}\sum_{n=1}^{N}q_{n}s_{n-1,k^{\prime}}s_{n,k^{\prime}}+\sum_{k=1}^{K_{x}}\sum_{n=1}^{N}q_{n}r_{n-1,k}r_{n,k}+Ls_{\mathrm{tot}}r_{\mathrm{tot}}$ is the total number of parameters.
The boosting-based hyperparameter selection procedure consists of two main stages: action order selection and rank reduction, which are detailed below. The overall procedure is summarized in Algorithm 2. Finally, the lag length $L$ is determined via a simple grid search to minimize the BIC over a predetermined candidate set $\{1,2,\ldots,L_{\max}\}$.
4.2.1 Action order selection
To construct the boosting-based algorithm, we first introduce the key notations and setup. Define $\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k})=\boldsymbol{\Lambda}_{x}(\{\alpha_{k}\})$, where the operator $\boldsymbol{\Lambda}_{x}(\cdot)$ is defined in (2.10) and $\boldsymbol{r}_{k}$ is the associated rank. It can be verified that for any two positive integer vectors $\boldsymbol{r}_{k,1}$ and $\boldsymbol{r}_{k,2}$, there exists a matrix $\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,3})$ with $\boldsymbol{r}_{k,3}\leq\boldsymbol{r}_{k,1}+\boldsymbol{r}_{k,2}$ by elementwise comparison, satisfying
$\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,3})=\boldsymbol{O}\left[\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,1}),\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,2})\right],$ (4.4)
where $\boldsymbol{O}$ is an invertible matrix that can be absorbed into $\boldsymbol{\Theta}$. The same applies to $\widetilde{\boldsymbol{\Lambda}}_{y}(\cdot)$ by defining $\widetilde{\boldsymbol{\Lambda}}_{y}(\beta_{k^{\prime}},\boldsymbol{s}_{k^{\prime}})=\boldsymbol{\Lambda}_{y}(\{\beta_{k^{\prime}}\})$, where $\boldsymbol{s}_{k^{\prime}}$ is the associated rank and $\boldsymbol{\Lambda}_{y}(\cdot)$ is defined similarly to (2.10). The proof is provided in Supplementary Material.
We refer to any model defined by (4.1) with $\boldsymbol{\Lambda}_{y}=\widetilde{\boldsymbol{\Lambda}}_{y}(\beta_{k^{\prime}},\boldsymbol{1}_{N})$ and $\boldsymbol{\Lambda}_{x}=\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{1}_{N})$ as a weak learner, denoted by $WL(\beta_{k^{\prime}},\alpha_{k})$, where $\boldsymbol{1}_{N}\in\mathbb{R}^{N}$ is a vector of ones. This weak learner serves as the base component in our boosting framework. The additive property introduced in (4.4) implies that the proposed model is additive with respect to rank increments: increasing the rank of a selected action order is equivalent to concatenating additional weak learners that share the same action order.
To reduce computational complexity, we predefine candidate sets of action orders, denoted by $\mathcal{P}_{y}^{(0)}$ and $\mathcal{P}_{x}^{(0)}$, based on prior knowledge of the data. These sets contain all feasible action orders to be considered for responses and predictors, respectively. For notational convenience, we abuse $K_{y}$ and $K_{x}$ to denote the number of action orders in $\mathcal{P}_{y}^{(0)}$ and $\mathcal{P}_{x}^{(0)}$ in this section.
To track how many times each action order is selected during the boosting process, we define the evolving action sets at iteration $v$ as $\mathcal{P}_{y}^{(v)}=\{\beta_{1}(s_{1}^{(v)}),\ldots,\beta_{K_{y}}(s_{K_{y}}^{(v)})\}$ and $\mathcal{P}_{x}^{(v)}=\{\alpha_{1}(r_{1}^{(v)}),\ldots,\alpha_{K_{x}}(r_{K_{x}}^{(v)})\}$, where $s_{k^{\prime}}^{(v)}$ and $r_{k}^{(v)}$ indicate the number of times action orders $\beta_{k^{\prime}}$ and $\alpha_{k}$ have been selected up to iteration $v$. We set $\mathcal{P}_{y}^{(0)}=\{\beta_{1}(0),\ldots,\beta_{K_{y}}(0)\}$ and $\mathcal{P}_{x}^{(0)}=\{\alpha_{1}(0),\ldots,\alpha_{K_{x}}(0)\}$ for initialization. For each selected action order $\beta_{k^{\prime}}(s_{k^{\prime}}^{(v)})\in\mathcal{P}_{y}^{(v)}$ and $\alpha_{k}(r_{k}^{(v)})\in\mathcal{P}_{x}^{(v)}$, we let $\boldsymbol{s}_{k^{\prime}}^{(v)}=s_{k^{\prime}}^{(v)}\boldsymbol{1}_{N}$ and $\boldsymbol{r}_{k}^{(v)}=r_{k}^{(v)}\boldsymbol{1}_{N}$. These sets and ranks are then used in Algorithm 1 to estimate $\boldsymbol{\Lambda}_{y}^{(v)}$, $\boldsymbol{\Lambda}_{x}^{(v)}$ and $\boldsymbol{\Theta}^{(v)}$, yielding the prediction $\widehat{\boldsymbol{y}}_{t}^{(v)}$.
[…]
$s_{\tilde{k}^{\prime}}^{(v+1)}=s_{\tilde{k}^{\prime}}^{(v)}+1,\quad r_{\tilde{k}}^{(v+1)}=r_{\tilde{k}}^{(v)}+1$
while keeping all other counts unchanged. By iteratively adding weak learners structured according to selected action orders, the model progressively refines its representational capacity in a data-adaptive and interpretable way.
4.2.2 Rank reduction
Algorithm 2 Boosting-Based Algorithm for Hyperparameter Selection
Input: Data $\{(\boldsymbol{y}_{t},\boldsymbol{x}_{t})\}_{t=1}^{T}$, initial sets $\mathcal{P}_{y}^{(0)}$ and $\mathcal{P}_{x}^{(0)}$, and maximum iterations $V_{\max}$.
Initialize: Set $\widehat{\boldsymbol{y}}_{t}^{(0)}=\boldsymbol{0}$ for all $t$ and $v=0$.
repeat
  Evaluate all candidate action order pairs $(\beta_{k^{\prime}},\alpha_{k})\in\{\beta_{1},\ldots,\beta_{K_{y}^{0}}\}\times\{\alpha_{1},\ldots,\alpha_{K_{x}^{0}}\}$:
  for all $(\beta_{k^{\prime}},\alpha_{k})$ do
   Fit weak learner $WL(\beta_{k^{\prime}},\alpha_{k})$ to regress $\boldsymbol{y}_{t}-\widehat{\boldsymbol{y}}_{t}^{(v)}$ on $\boldsymbol{x}_{t}$.
  end for
[…]
  Update action counts: $s_{k^{\prime}}^{(v+1)}=s_{k^{\prime}}^{(v)}+\mathbb{I}\{k^{\prime}=\tilde{k}^{\prime}\}$, $r_{k}^{(v+1)}=r_{k}^{(v)}+\mathbb{I}\{k=\tilde{k}\}$
  Update action sets: $\mathcal{P}_{y}^{(v+1)}=\{\beta_{k^{\prime}}(s_{k^{\prime}}^{(v+1)})\}_{k^{\prime}=1}^{K_{y}}$, $\mathcal{P}_{x}^{(v+1)}=\{\alpha_{k}(r_{k}^{(v+1)})\}_{k=1}^{K_{x}}$
  Refit the full model using Algorithm 1 with $\mathcal{P}_{y}^{(v+1)}$ and $\mathcal{P}_{x}^{(v+1)}$ to obtain $\widehat{\boldsymbol{y}}_{t}^{(v+1)}$.
  $v\leftarrow v+1$
until BIC increases or $v=V_{\max}$
 Perform rank reduction: $(\mathcal{P}_{y},\mathcal{P}_{x})=\text{RR}(\mathcal{P}_{y}^{(v)},\mathcal{P}_{x}^{(v)})$
Return: $\mbox{$\mathscr{A}$}]_{N}$ by applying Algorithm 1 with action sets $\mathcal{P}_{y}$ and $\mathcal{P}_{x}$
According to Proposition 1, multiple combinations of action order sets $(\mathcal{P}_{y}^{(v)},\mathcal{P}_{x}^{(v)})$ may yield equivalent representations of the same model. Among them, we prefer the representation that achieves the lowest model complexity, as measured by BIC defined in (4.3). To this end, we introduce a procedure for selecting the optimal pair $(\mathcal{P}_{y},\mathcal{P}_{x})$ after the action order selection stage. We refer to this procedure as the rank reduction step, denoted by $\text{RR}(\cdot)$.
The rank reduction procedure is applied to both $\mathcal{P}_{x}$ and $\mathcal{P}_{y}$. We illustrate the process using $\mathcal{P}_{x}$ as an example. The procedure iterates sequentially over action orders $\alpha_{k}$ with $k=1,\ldots,K_{x}$. For a given action order $\alpha_{k}$ with associated rank $\boldsymbol{r}_{k}=(r_{1,k},\ldots,r_{N,k})$, we apply Proposition 1 to search for an equivalent but more efficient representation by exploring all alternative action orders within $\mathcal{P}_{x}$.
[…]
An identical procedure is then applied to $\mathcal{P}_{y}$, ensuring both $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$ are represented in their most parsimonious forms. The final output of the rank reduction step is the optimal pair $(\mathcal{P}_{y},\mathcal{P}_{x})$ corresponding to the most efficient, yet equivalent, representation of the model structure. We summarize the overall hyperparameter selection procedure into Algorithm 2.
5 Simulation studies
This section conducts three simulation studies to evaluate the finite-sample performance of our OLS estimator proposed in Algorithm 1, investigate the effects of action order misspecification, as discussed in Proposition 1, and compare the performance of different feature extraction methods for hierarchical data under the framework of supervised factor modeling.
In the first experiment, the data generating process follows (2.12) with $L=2$ and $N=3$ and equal dimension sizes, i.e., $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{3}=([\mbox{$\mathscr{A}$}_{1}]_{3},[\mbox{$\mathscr{A}$}_{2}]_{3})$ and $\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{q\times q\times q}$. For simplicity, let intrinsic ranks $r_{n,1}=r_{n,2}=r$, $s_{n,1}=s_{n,2}=s$, where $r=s+1$. We consider three types of sub-Gaussian noise: (i) i.i.d. uniform noise, (ii) i.i.d. standard normal noise, and (iii) correlated gaussian noise. The detailed data generation process is relegated to Supplementary Material.
From Theorem 2, the estimation error converges at a rate of $\sqrt{d_{\mathrm{AR}}/T}$, where $d_{\mathrm{AR}}=2(K_{x}r^{2}+K_{y}s^{2})Nq+2LK_{x}K_{y}rs$. To numerically verify this convergence rate, we consider three settings: (a) Fix $(r,T)=(3,2500)$ and vary $q\in\{10,11,12,13,14\}$; (b) Fix $(q,T)=(10,2500)$ and vary $r\in\{3,4,5,6,7\}$; (c) Fix $(q,r)=(10,3)$ and vary $T\in\{833,1000,1250,1670,2500\}$. In all cases, we apply Algorithm 1 to estimate $\mbox{$\mathscr{A}$}_{\mathrm{AR}}$. Figure 2 presents the estimation error $\|\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{LR}}-\mbox{$\mathscr{A}$}^{*}_{\mathrm{LR}}\|_{\mathrm{F}}$ for these settings, averaged over 200 replications. The observed linearity in Figure 2 confirms the theoretical convergence rate in Theorem 2. It verifies that the estimation error depends on $q$ linearly, and $r$ quadratically, thereby demonstrating the efficiency of our proposed method.
[…]
In the second experiment, we investigate the impact of action order misspecification, as discussed in Proposition 1. We set $M=3$, then the action set is $\mathcal{P}(3)$, which consists of all permutations of $\{1,2,3\}$ with cardinality $|\mathcal{P}(3)|=6$. The data are generated by $\boldsymbol{y}_{t}=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha_{1});\{\boldsymbol{G}_{m,1}\}_{m=1}^{3})$, where the true action order is $\alpha_{1}=123$. For simplicity, let $r_{1,k}=r_{2,k}=r_{3,k}=r_{k}$ for $k=1,\ldots,6$, and $\boldsymbol{G}_{m,1}$’s are generated similarly as in the first experiment. We consider the i.i.d. normal setting for $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{r_{k}\times r_{k}\times r_{k}}$ with $\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\sim N(\boldsymbol{0},\boldsymbol{C})$, where $\boldsymbol{C}$ is a random matrix.
By Proposition 1, for each $\alpha_{k}\in\mathcal{P}(3)$, there exists a corresponding set of $\{\boldsymbol{G}_{m,k}\}_{m=1}^{3}$ such that $\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha_{1});\{\boldsymbol{G}_{m,1}\}_{m=1}^{3})=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha_{k});\{\boldsymbol{G}_{m,k}\}_{m=1}^{3})$, with the cost of potentially larger intrinsic rank $r_{k}$. We apply Algorithm 1 to estimate $\boldsymbol{G}_{m,k}$’s for each $k$, and compute the mean squared error (MSE) of the predictions. Figure 3(a) presents the MSEs for different values of $r_{k}$’s averaged over 50 replications. For both settings of $\mbox{$\mathscr{X}$}_{t}$, we observe that when $r_{k}$ is small, the MSEs for the misspecified action orders are significantly larger than that of the true action order $\alpha_{1}$. However, as $r_{k}$ increases, MSEs for all action orders converge to the same level as that of the true action order $\alpha_{1}$. This aligns with our theoretical findings in Proposition 1.
Figure 3: (a) Average MSEs under different action orders with varying intrinsic rank $r_{k}$. (b) Average rolling forecast MSEs for HOFM, HFM, and our proposed method under supervised factor modeling framework.
In the last experiment, we compare our proposed method with two hierarchical feature extraction methods: HFM (schmid1957development) and HOFM (yung1999relationship), under the framework of SFM. The data generating process follows (2.12) with the coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}$ constructed by both HOFM and our proposed method. Details of the data generation and training procedure are relegated to Supplementary Material.
To compare the one-step ahead forecast performance of HOFM, HFM, and our proposed method, we apply Algorithm 1 to estimate the coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}$ for our model, and use the alternating least squares algorithm to estimate the coefficient tensor for HFM and HOFM. Figure 3(b) presents the average rolling forecast MSEs over 50 replications. We observe that our method consistently outperforms both HFM and HOFM, regardless of whether the data are generated by HOFM or our method. This demonstrating the efficiency of our model in capturing hierarchical structures with shared components within each mode in loading matrices for feature extractions.
6 Real data analysis
In this section, we demonstrate the application of our proposed model to the International Personality Item Pool (IPIP) NEO-120 personality inventory (johnson2014measuring), using data collected from June 1, 2015 to April 30, 2025. The dataset is available in the Hugging Face Datasets repository: https://huggingface.co/datasets/ecorbari/IPIP120-SCORES. It includes scores for the five major personality traits (Neuroticism, Extraversion, Conscientiousness, Agreeableness, and Openness to Experience), assessed by 24 questions. Participants are categorized by gender (two groups) and age (six groups: 10-19, 20-29, 30-39, 40-49, 50-59, and 60+), resulting in 12 demographic groups. To analyze trends in personality traits over time, we compute monthly average scores for each question (recorded on a 1-5 Likert scale) within each demographic group, forming a tensor time series $\{\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{12\times 5\times 24}\}_{t=1}^{T}$ with $T=115$ observations. There is one missing entry in the dataset, which we impute using spline interpolation. For preprocessing, the data are first detrended by first-order differencing, and then standardized to have zero means.
The tensor autoregression model (2.12) is employed with each coefficient tensor $\mbox{$\mathscr{A}$}_{l}\in\mathbb{R}^{12\times 5\times 24\times 12\times 5\times 24}$. Observations are divided into a training set with the first 103 months and a testing set with the last 12 months. To enable a feasible estimation, we consider our proposed model, SFM based on CP and TT decompositions, and HOFM under SFM frameworks for feature extraction with the lag order fixed at $L=1$ by jointly consider the PACFs of all entries in data. We omit the comparison with Tucker decomposition and HFM based on SFM frameworks due to thier large number of parameters in the coefficient matrix, which makes them infeasible for this data analysis. For our proposed model, from prior knowledge of the data, we set the candidate action order sets as $\mathcal{P}_{x}^{(0)}=\mathcal{P}_{y}^{(0)}=\{(3,2,1),(1,3,2)\}$, i.e., the features are extracted by the order of questions, personalities, and demographic groups, or the order of demographic groups, questions, and personalities. We then apply Algorithm 2 to get the selected intrinsic ranks $\boldsymbol{r}_{1}=(1,1,2)$, $\boldsymbol{r}_{2}=(9,1,2)$, $\boldsymbol{s}_{1}=(3,1,2)$, and $\boldsymbol{s}_{2}=(3,2,2)$. For the two tensor decomposition based SFMs, we adopt the ALS methods in zhou2013tensor for CP-based tensor autoregression, and wang2022high for Tucker-based tensor autoregression. For TT-based SFM, we follow si2024efficient and qin2025computational to perform Riemannian gradient descent for estimation. The intrinsic ranks for these methods are selected by their respective methods.
We next evaluate the rolling forecast performance on the testing set of these four methods, i.e., fit the model to historical data up to month $t$ to forecast the data at month $t+1$ for $t=103,\ldots,114$. Specifically, we use mean squared forecast errors (MSFE) and mean absolute forecast errors (MAFE) to assess the one-step ahead forecast accuracy. Table 2 summarizes the results, where our proposed method outperforms all competing methods in both MSFE and MAFE metrics. This result supports the existence of hierarchical structures in the IPIP NEO-120 data, which can be effectively captured by our proposed model. Moreover, the superior performance of our method over TT decomposition based SFM highlights the advantage of our model by considering multiple action orders for general hierarchical data. And the better performance compared to HOFM based SFM demonstrates the benefit of improved efficiency in feature extraction by sharing components within each mode in loading matrices. In conclusion, our proposed model provides an efficient and interpretable approach for modeling tensor time series data with hierarchical structures.
Table 2: Mean squared forecast errors (MSFE) and mean absolute forecast errors (MAFE) of our model and three competing methods on the testing set of the IPIP NEO-120 data. The smallest numbers in each row are in bold.
Ours CP $T$ HOFM
MSFE 138.78 142.28 145.73 142.15
MSAE 279.28 280.63 283.43 280.58
Finally, we zoom in on the estimated coefficient tensor from our proposed model to interpret the hierarchical structures in the IPIP NEO-120 data and the importance of different action orders. We discuss the component matrices involved in feature extraction for responses under action order $(1,3,2)$ as an example. Figure 4(a) presents the heatmap of the component matrix $\boldsymbol{H}_{1,1}^{\top}$, which correspond to the first step of feature extraction, summarizing the information across the 24 questions for each personality trait. In addition, Figure 4(b) displays $\boldsymbol{H}_{2,1}^{\top}$ that represents the combined effects of personality traits and questions by a single extracted feature. Lastly, Figure 4(c) illustrates $\boldsymbol{H}_{3,1}^{\top}$, capturing the interactions between demographic groups and the combined features from questions and personality traits. The response tensor $\mbox{$\mathscr{Y}$}_{t}$ is then summarized into two factors as its low-dimensional representation. Component matrices with the second action order $(1,3,2)$ for responses and those for predictors with both action orders can be found in Supplementary Material, which have similar interpretations as those for predictors. Finally, Figure 4(d) presents the heatmap of $\boldsymbol{\Theta}$, the estimated coefficient matrix that captures the relationships between the extracted factors from predictors and responses by two action orders. We observe that features extracted by action order $(3,2,1)$ have a more significant impact on extracted features from responses under both action orders compared to those extracted by action order $(1,3,2)$, as indicated by the larger absolute values in the corresponding blocks of $\boldsymbol{\Theta}$.
Figure 4: Heatmaps of (a) $\boldsymbol{H}_{1,1}^{\top}$, (b) $\boldsymbol{H}_{2,1}^{\top}$, (c) $\boldsymbol{H}_{3,1}^{\top}$, and (d) $\boldsymbol{\Theta}$.
7 Conclusion and discussion
In this paper, we introduce a novel autoregressive model under SFM framework for tensor time series data that effectively captures hierarchical structures through the incorporation of multiple action orders in the construction of the coefficient tensor. We employ a structure of shared components within each mode in loading matrices for efficient feature extraction. An alternating least squares algorithm is developed for parameter estimation, and a boosting-based method is proposed for hyperparameter selection, which aims to reveal the underlying hierarchical structures present in the data. Theoretical analyses demonstrate the consistency of our estimator under sub-Gaussian noise with misspecification of action orders. Simulation studies and real data analysis illustrate the superior performance of our method compared to existing approaches.
For future research, we may extend the proposed model to accommodate more complex time series structures, such as moving average (MA) or ARMA models. We can also develop robust estimation techniques for handling heavy-tailed noise distributions, so that our methods can be deployed to a wider range of real-world applications. Additionally, investigating statistical inference procedures for the selected action orders and intrinsic ranks would provide valuable insights into the interpretability and reliability of the model in practical settings.